# Local-Cloud Inference Offloading for LLMs in Multi-Modal, Multi-Task, Multi-Dialogue Settings

### v1

@@ III-D Uncertain Response Score Estimation
Response Quality Evaluation. The response score $S_{R_{t}}$ measures the response quality generated by the selected LLM in each dialogue. In multi-task scenarios, evaluating the response quality goes beyond merely verifying control actions within an assistant system; it critically involves assessing the actual utility of responses in aiding human users, particularly in query tasks. We utilize the quantitative evaluation scheme outlined by [40, 41, 42], as illustrated in Fig. 4. This involves using two additional text-only LLMs specifically to assess the quality of candidate LLM services. The reference LLM processes a prompt and a reference response to produce a polished reference response, serving as a benchmark for evaluation. The judge LLM then evaluates both the candidate’s response and the reference response, scoring them based on helpfulness, relevance, accuracy, and detail to derive the response score $S_{R_{t}}$, which is calculated as the ratio of the candidate’s LLM score to the reference score.
Uncertain Response Score Estimation. Integrating RCRL directly into real-time LLM inference is impractical, especially in our scenario where four LLMs operate in tandem: local and cloud-based LLMs generate responses, while reference and judge LLMs evaluate their quality. Consequently, we first archive the LLM inference processes into a dataset for subsequent offline RCRL. However, this approach introduces two main forms of uncertainty: (i) non-deterministic evaluation (NDE), in which identical state-action pairs can yield different response quality scores; and (ii) the potential presence of out-of-distribution (OOD) state-action pairs in the dataset. In contrast to [43], which aims to avoid selecting OOD actions, our goal is to more accurately estimate response quality under OOD actions rather than constrain them. This is because the states and actions in our M4A1 dataset are randomly generated (e.g., the current task, selected LLM, and selected modality), making it inevitable that some state-action pairs will not appear in the dataset. To address these uncertainties, we estimate the response score $S_{R}^{\prime}$ for a pending state-action pair $(s^{\prime},a^{\prime})$ by leveraging its nearest neighbors in the dataset. Specifically, let $\{(s_{i},a_{i})\}_{i=1}^{k}$ be the $k$ state-action pairs nearest to $(s^{\prime},a^{\prime})$ in Euclidean distance, and let $\{S_{R_{(i)}}\}$ denote their known response scores. We then estimate $S_{R}^{\prime}$ as a weighted average:
@@
$S_{R}^{\prime}=\frac{\sum_{i=1}^{k}\mathrm{w}_{i}\cdot{S_{R}}_{(i)}}{\sum_{i=1}^{k}\mathrm{w}_{i}}\quad\mathrm{s.t.}\quad\min_{(s_{i},a_{i})}\left\|(s^{\prime},a^{\prime})-(s_{i},a_{i})\right\|,$ (12)
@@
where $\mathrm{w}_{i}=\frac{1}{\|(s^{\prime},a^{\prime})-(s_{i},a_{i})\|}$ represents the weight assigned to each known state-action pair $(s_{i},a_{i})$, inversely proportional to their distance from $(s^{\prime},a^{\prime})$. In this way, we achieve a more robust offline estimation of response quality, effectively handling both NDE and OOD uncertainties within RCRL for LLM inference.
III-E Association of Task and Multi-modal Data Source
@@
LCIO is designed to operate in multi-modal, multi-task scenarios. In such settings, it is important to quantify the relevance between user’s text prompts and multi-modal data sources. This helps the user select the most pertinent data modalities to offload to the cloud LLM along with the text prompt for inference. To facilitate this, we introduce an association metric to gauge the relevance of each data modality to the content of the text prompt. Consequently, we employ feature extractors to transform both tasks and modalities into feature vectors of the same dimensionality, which are then used to compute their similarity. Given this premise, we propose leveraging pre-trained CLIP model [44], a transformer-based model comprising a text encoder and an image encoder, to extract features from prompts and multi-model data sources. The pre-trained CLIP has been extensively trained on a vast array of image-text pair data, effectively linking visual concepts with text. In contrast to traditional image classifiers trained on fixed datasets like ImageNet, CLIP can handle both text and images, and has been widely proven to be applicable to nearly any visual classification task. We employ a normalized cross-modality similarity metric to calculate the association score between the text prompt $P$ and the data modality $m$, which is defined as:
@@
$\Lambda\left(P,D^{(m)}\right)=\frac{\left\langle\omega_{T}(P),\omega_{I}(D^{(m)})\right\rangle}{\left\|\omega_{T}(P)\right\|\cdot\left\|\omega_{I}(D^{(m)})\right\|},$ (13)
@@
where $\omega_{T}:P\rightarrow\mathbb{R}^{d}$ and $\omega_{I}:D^{(m)}\rightarrow\mathbb{R}^{d}$ represent the text and image encoder functions, respectively, mapping inputs to a shared $d$-dimensional embedding space. The denominator ensures normalization through $L_{2}$ norms, producing similarity scores bounded in $[-1,1]$.
III-F Latency Model
@@
Local Computational Latency $\psi_{\text{Local}}$. The computational latency of the local LLM, or inference time, is influenced not only by the length of the prompt $|P|$ and the response $|R|$, but also significantly by the number of LLM parameters. Following [45, 46], we assume that the computational latency is proportional to the number of parameters, though other latency models can be easily applied. For our local LLM, comprising $|\text{LLM}_{\text{Local}}|$ parameters and requiring $2|\text{LLM}_{\text{Local}}|$ floating point operations for forward propagation, the computational demand for each token in a basic processing length $|P|_{\text{base}}$ is calculated as $\frac{2|\text{LLM}_{\text{Local}}|}{|P|_{\text{base}}}$ floating-point operations. Additionally, other functions may also describe the relationship between the number of parameters and latency, reflecting the complexity of the model and the efficiency of the hardware. The latency calculation considers the theoretical peak performance of the local device’s GPU, $TF_{\text{peak}}$ TFLOPS, leading to the computational latency being defined as:
$\psi_{\text{Local}}=\frac{2|\text{LLM}_{\text{Local}}|(|P|+|R|)}{|P|_{\text{base}}\times TF_{\text{peak}}}.$ (14)
@@
Cloud Interaction Latency $\psi_{\text{Cloud}}$. The interaction time with the cloud LLM encompasses the entire process from sending the prompt to receiving the response. This includes uploading the prompt and modalities, inference by the cloud LLM, and downloading the response. We record the interaction time with the cloud LLM $\psi_{\text{Cloud}}$, for different sets of uploaded modalities $\mathcal{M}$, as described in Section IV-A.
Algorithm 2 M4A1 Dataset Collection
Input: Local LLM, cloud LLM, reference LLM, judge LLM, image dataset (e.g., ActionSense), instruction dataset, time span $(\tau)$
Output: M4A1 dataset
1:while collecting data do
2:    Randomly select a moment in the scenario and retrieve images with various modalities.
3:    for a random number of dialogues from 2 to $\tau$ do
4:         Randomly select a user prompt from the instruction dataset.
5:         Randomly choose either the local LLM or the cloud LLM.
6:         if the chosen LLM is the cloud LLM then
7:             Randomly select which modality (text-only, a single view, or multiple views) to upload to the cloud LLM.
8:         end if
9:         Generate a response using the selected LLM based on the context of prior dialogues, the current prompt, and the selected modality.
10$\triangleright$ Eqs. (1), (2)
11:         The reference LLM generates a reference response based on the prompt and a human-generated reference response.
12:         The judge LLM assigns a response score given the prompt, the candidate response, and the reference response.
13:         Compute the local latency $\psi_{a}$ or record the real interaction time with the cloud LLM provider. $\triangleright$ Eq. (14)
14:         Compute the usage cost $\phi_{a}$ for either the local or cloud LLM. $\triangleright$ Eqs. (15), (16)
15:    end for
16:    for each prompt $P$ and each multimodal data $D^{(m)}$ in the current sample do
17:         Compute the association $\Lambda(P,D^{(m)})$ by the pre-trained CLIP model. $\triangleright$ Eq. (13)
18:    end for
19:end while
III-G Usage Cost Model
@@
During the inference process, LLMs incur computational costs. For the local LLM, it manifests as energy consumption, while for the cloud LLM, it translates into the service fee paid to the providers. The costs associated with running LLMs are determined by factors such as computational complexity, processing duration, and computing equipment.
@@
Local LLM Usage Cost $\phi_{\text{Local}}$. The energy consumption of the local LLM is influenced by inference time and the power consumption of the computing hardware, which are in turn affected by the local device’s computational capacity, the sizes of the prompt and response, and the size of the LLM, all encapsulated within the inference time $\psi_{\text{Local}}$. For a given local device hardware, the maximum power consumption is $W_{\text{max}}$ watts. Furthermore, by utilizing energy costs $\kappa$ (such as local electricity rates) and after appropriate unit conversion, the local LLM usage cost (in USD) is defined as:
$\phi_{\text{Local}}=\psi_{\text{Local}}\times W_{\text{max}}\times\kappa.$ (15)
@@
Cloud LLM Usage Function $\phi_{\text{Cloud}}$. For cloud-based LLMs, in addition to the inference costs of prompts and responses, the inference costs of data modalities must also be considered. These costs are typically available on the service provider’s website. Taking GPT-4o as an example, the prompt cost rate is $\varphi_{P}=\$0.005/1K$ tokens, the response cost rate is $\varphi_{R}=\$0.015/1K$ tokens, and the data modality costs are depending on size and resolution. Therefore, the usage cost of the cloud LLM can be defined as follows:
$\phi_{\text{Cloud}}=\varphi_{P}\left(|P|\right)+\varphi_{R}\left(|R|\right)+\sum_{m\in\mathcal{M}}\varphi_{m}\left(\big{|}D^{(m)}\big{|}\right).$ (16)
Figure 4: Illustration of M4A1 Dataset. Multi-modal data include first-person view, side view, and overhead view. Client’s prompt dialogues represent four tasks, labeled as Assistant, Recommendation, Query, and Message Editing. LLMs’ dialogues originate from different LLMs, specifically Local LLM and Cloud LLM. According to the evaluation by the judge LLM, Response Score is used to quantify the response quality of the candidate LLMs.
IV Experimental Results
IV-A M4A1 Dataset Construction
IV-A1 Overview of Proposed M4A1
@@
To evaluate performance in practical settings, we develop M4A1, the first multi-modal, multi-task, multi-dialogue, and multi-LLM dataset. The data collection method is detailed in Alg. 2, and an overview of M4A1 is depicted in Fig. 4. M4A1 consists of 6000 samples, each meticulously compiled by randomly and sequentially choosing tasks, prompts, modalities, and candidate LLMs to ensure comprehensive coverage across different configurations. The key characteristics of M4A1 can be summarized as follows:
@@
  • Four LLMs contain a local LLM (Phi-3-mini [34]), a cloud LLM (GPT-4o [18]), a reference LLM (GPT-4o), and a judge LLM (GPT-4o).
@@
  • Four Metrics include response score (Sec. III-D), association (Sec. III-E), latency (Sec. III-F), and usage cost (Sec. III-G).
IV-A2 M4A1 Construction Details
To construct the M4A1 dataset, we integrate two predetermined datasets that provide complementary modalities, images and text. First, we use the ActionSense dataset [47] to provide images, offering various camera views of human along with labels for each kitchen activity within the scenario, as illustrated in the bottom left of Fig. 4. ActionSense is a large dataset of human kitchen activities that includes multi-modal wearable sensor data and environmental camera data, capturing 20 different kitchen activities such as peeling cucumbers, slicing potatoes, and cleaning plates with a sponge. These annotated kitchen activity labels function as responses to queries such as “What activity am I doing right now?” Given that current LLMs rarely process time-series sensor data (such as data from gyroscopes and accelerometers), we consider three types of image modalities: first-person images from a wearable sensor, and side and overhead views from environmental cameras. Second, the instruction dataset is used to provide structured text data, enriching the dataset with linguistic elements crucial for LLM inference and evaluation. This dataset provides text for four different tasks with prompts and reference responses. Each task contains five unique prompts, with examples in Fig. 5. For each prompt, a reference response is also provided. Note that the reference responses do not affect the evaluation because the reference LLM will reprocess the content and format of the reference responses.
Each sample in M4A1 is constructed based on a set of three multi-modal data from ActionSense, paired with multiple randomly selected prompt and reference response pairs from the instruction dataset.
I’m feeling a little hot right now, please turn the temperature down a bit. Please play some Christmas music. Please turn on all the lights in the kitchen.
Please recommend some music for me to play right now. Please recommend two dishes that I can cook based on what I have. What is the next recommended step for finishing the meal I am cooking?
What activity am I doing right now? Where did I put my dishes? What are the calorie counts for potatoes and cucumbers?
Please draft a text message asking my friend to bring some ketchup… Please draft a memo to remind me to buy some eggs tomorrow. Please send a message to my home that I’m stuck in traffic and will be home in…
Figure 5: Exemplary Samples in the Instruction Dataset. The dataset consists of prompts that fall into one of the following task categories: Assistant, Recommendation, Query, and Message Editing.
IV-B Experiment Setup
@@
Training Setup. The M4A1 dataset is split into a training set and a test set in an 8:2 ratio. The training set is used to train RLs while the test set is used to evaluate the performance at inference. When implementing our approach, we consider several popular discrete RL algorithms, including PPO [36], DQN [37], and A2C [35]. We consider these methods with and without resource constraints (corresponding to cases with $\lambda>0$ and $\lambda=0$ in Eq. (11)), and denote RC-PPO, RC-DQN, and RC-A2C as the methods with the resource constraints. The policies are parameterized by multilayer perceptrons (MLP), with a total of 30,000 time steps set for training. Default parameters are used as specified in the Stable Baselines3 library [48]. For each experiment, we repeat the process three times and report the standard deviation as the error bar. We implement the LCIO system and the baselines on PyTorch and conducted the experiments on an NVIDIA A100 GPU with 40 GB of memory.
TABLE I: Main Result - LCIO with RC-A2C achieves a balance between local and cloud LLMs while adaptively selecting modalities. It outperforms all baselines with superior response scores, reduced latency, and lower usage costs, without any constraint violations.
@@
Method Response Latency Usage Cost Overall Local Cloud - Num. Selected Modalities 1 Constraint Violation 2 ($\downarrow$)
@@
LLM-as-Agent for Inference Offloading (Ignored LLM Agent’s Latency and Usage Cost)
@@
Comparison with SOTA
AIwRG [33] 1.02 $\pm$0.05 23.73 $\pm$0.98 44.01 $\pm$3.50 0.69 $\pm$0.06 113 244 0 0 3852 4.68 $\pm$0.18 9.26 $\pm$1.83
PerLLM [49] 0.97 $\pm$0.06 21.12 $\pm$0.08 1.60 $\pm$0.01 0.66 $\pm$0.06 281 3922 0 0 0 0.00 $\pm$0.00 0.00 $\pm$0.00
LCIO (PPO) 1.02 $\pm$0.00 17.89 $\pm$0.82 22.48 $\pm$1.48 0.76 $\pm$0.01 142 40 2499 1355 121 0.70 $\pm$0.77 1.66 $\pm$2.35
LCIO (DQN) 1.02 $\pm$0.08 19.57 $\pm$2.18 26.23 $\pm$6.36 0.75 $\pm$0.05 0 5 1820 2367 0 0.81 $\pm$0.57 0.00 $\pm$0.00
LCIO (A2C) 1.05 $\pm$0.03 18.81 $\pm$3.44 26.44 $\pm$13.05 0.79 $\pm$0.07 98 0 2753 210 1168 1.48 $\pm$2.09 2.38 $\pm$3.36
LCIO (RC-PPO) 1.03 $\pm$0.03 18.33 $\pm$1.87 22.43 $\pm$3.57 0.77 $\pm$0.04 90 105 2599 1309 110 0.45 $\pm$0.64 0.00 $\pm$0.00
LCIO (RC-DQN) 1.05 $\pm$0.05 18.87 $\pm$0.81 23.29 $\pm$3.59 0.78 $\pm$0.06 46 174 2249 1636 85 0.53 $\pm$0.33 0.42 $\pm$0.60
LCIO (RC-A2C) 1.09 $\pm$0.08 16.63 $\pm$0.67 17.97 $\pm$1.18 0.85 $\pm$0.07 74 12 3871 225 6 0.00 $\pm$0.00 0.00 $\pm$0.00
Ours LCIO System - Ablation Study (Using RC-A2C as Backbone)
w/o LLM Sel. 3 0.85 $\pm$0.00 9.25 $\pm$1.02 11.67 $\pm$3.14 0.72 $\pm$0.02 2128 6 1228 830 1 0.00 $\pm$0.00 0.00 $\pm$0.00
@@
1 Num. Selected Modalities refers to the number of modalities selected in a single dialogue. For example, with two selected modalities, combinations include {first-person, side}, {first-person, overhead}, {side, overhead}.  2 Constraint Violation refers to the instances where the cumulative resource consumption caused by multiple consecutive actions in each sample exceeds the threshold value, as defined by $\max\left\{g_{j}(\theta),0\right\}$ in Eq. (11).  3 w/o LLM Sel., w/o Modality Sel., and w/o Score Est. refer to without LLM selection, without modality selection, and without uncertain response score estimation, respectively.
@@
Baselines. We compare the RL algorithms with three naive baselines: random (with random LLM and modality selection), local (using only the local LLM), and cloud (using only the cloud LLM, but with random modality selection). In addition, we replace RL with LLM-as-agent for LLM and modality selection. Specifically, the current state and possible actions are translated into natural language descriptions and provided as contextual input to the LLM agent for decision-making. These LLM agents do not access any information beyond what is provided to RL. The LLM agents include Phi-3-mini [34], Phi-3.5-mini [34], Llama-3.2-3B [6], Llama-3.1-8B [6], Mistral-7B-v0.3 [50], FLAN-T5-large [51], FLAN-T5-xl [51], Gemma-2-2B [3], Gemma-2-9B [3], GPT-3.5-turbo [52], GPT-4o-mini [53], GPT-4o [18], OpenAI o1-mini [54], OpenAI o1 [54], and OpenAI o3-mini [55]. We reiterate that none of the existing works can be directly applied to our setting, which features multi-modal, multi-task, and multi-dialogue characteristics. Therefore, we reimplement AIwRG [33], an active inference method, and PerLLM [49], a bandit-based approach using Upper Confidence Bound (UCB) for exploration-exploitation trade-off, integrating the settings introduced in this paper to adapt them to our research problem.
Default Settings. The default experimental settings are as follows unless otherwise stated; we also study the effects of various system parameters. Specifically, we use RC-A2C for the RL method, with $\alpha=1$, $\beta_{\Lambda}=1/3$, $\beta_{\psi}=1/3$, and $\beta_{\phi}=1/3$ for the reward function weights, a 30 second latency constraint, a 0.05 USD usage cost constraint, and Jetson TX2 as the device type. These settings serve as the baseline from which variations are systematically introduced to assess their effects. Additionally, we set the time span $\tau=5$, the penalty coefficient $\lambda=10$, $k=5$ nearest neighbors for estimating the response score, and $\kappa=4.63\times 10^{-8}$ (Joules per USD) based on the average electricity price in the US  [^2] https://www.energybot.com/electricity-rates/.
IV-C Main Experimental Results

### v3

@@ 3.4.1. Response Quality Evaluation
The response score $S_{R_{t}}$ measures the response quality generated by the selected LLM in each dialogue. In multi-task scenarios, evaluating the response quality requires assessing the actual utility of responses in aiding human users, particularly in query tasks. We employ the quantitative evaluation scheme outlined by (Liu et al., 2024), illustrated on the right side of Fig. 3. This involves using two additional text-only LLMs specifically to assess the quality of candidate LLM services:
  • The reference LLM processes a prompt and a reference response to produce a polished reference response, serving as a benchmark for evaluation.
  • The judge LLM then evaluates both the candidate’s response and the reference’s response, scoring them based on helpfulness, relevance, accuracy, and detail.
The response score $S_{R_{t}}$ for the dialogue $t$ is then calculated as the ratio of the candidate’s LLM score to the reference score.
3.4.2. Handling Uncertainty in Response Score Estimation
Integrating RCRL directly into real-time LLM inference is impractical, especially in our scenario where four LLMs operate in tandem: local and cloud-based LLMs generate responses, while reference and judge LLMs evaluate their quality. Consequently, we archive the LLM inference processes into a dataset for offline RCRL. However, this approach introduces two main forms of uncertainty:
  • (i)
    non-deterministic evaluation (NDE), in which identical state-action pairs can yield different response quality scores; and
  • (ii)
    the potential presence of out-of-distribution (OOD) state-action pairs in the dataset.
In contrast to (Ran et al., 2023), which aims to avoid selecting OOD actions, our goal is to more accurately estimate the response quality under OOD actions rather than constrain them. This adds flexibility for datasets where the states and actions are randomly generated (e.g., the current task, selected LLM, and selected modality), making it inevitable that some state-action pairs will not appear in the dataset (such as for our M4A1 dataset described in Sec. 4.1).
To address these uncertainties, we estimate the response score $S_{R}^{\prime}$ for a pending state-action pair $(s^{\prime},a^{\prime})$ by leveraging its nearest neighbors in the dataset. Specifically, let $\{(s^{i},a^{i})\}_{i=1}^{k}$ be the $k$ state-action pairs nearest to $(s^{\prime},a^{\prime})$ in Euclidean distance, and let $\{S_{R^{i}}\}_{i=1}^{k}$ denote their known response scores. We estimate $S_{R}^{\prime}$ as a weighted average over its neighbors:
@@ 3.4.2. Handling Uncertainty in Response Score Estimation
(11) $\small S_{R}^{\prime}=\frac{\sum_{i=1}^{k}{S_{R^{i}}/d_{i}}}{\sum_{i=1}^{k}1/d_{i}},\quad\mathrm{where}\quad d_{i}=\left\|[s^{\prime}\;a^{\prime}]-[s^{i}\;a^{i}]\right\|$
@@
represents the distance between the $k$ nearest state-action pairs and the pending state-action pair $(s^{\prime},a^{\prime})$, with [xy][x\;y] denoting vector concatenation. In this way, we achieve a more robust offline estimation of response quality, effectively handling both NDE and OOD uncertainties within RCRL for LLM inference.
3.5. Task-Modality Association
@@
TMO is designed to operate in multi-modal, multi-task scenarios. In such settings, quantifying the relevance between users’ text prompts and multi-modal data sources enables selection of the most pertinent multi-modal data sources to offload to the cloud LLM. Consequently, we employ feature extractors to transform both tasks and modalities into feature vectors of identical dimensionality for similarity computation. Given this premise, we leverage the pre-trained CLIP model (Radford et al., 2021), a transformer-based architecture with text and image encoders, to extract features from prompts and multi-model data sources. We employ a normalized cross-modality similarity metric to calculate the association score between a text prompt $P$ and data modality $m$, which is defined as:
@@
(12) $\small\Lambda\left(P,D^{(m)}\right)=\frac{\left\langle\omega_{T}(P),\omega_{I}(D^{(m)})\right\rangle}{\left\|\omega_{T}(P)\right\|\cdot\left\|\omega_{I}(D^{(m)})\right\|},$
@@
where $\omega_{T}:P\rightarrow\mathbb{R}^{d}$ and $\omega_{I}:D^{(m)}\rightarrow\mathbb{R}^{d}$ represent the text and image encoders, respectively, mapping inputs to a shared $d$-dimensional embedding space.
3.6. Latency Model
3.6.1. Local Computation Latency $\psi_{\text{Local}}$
@@
The computational latency of the local LLM, or say inference time, is influenced not only by the number of LLM parameters, but also significantly by the length of the prompt, denoted $|P|$, and of the response, $|R|$. Following (Canziani et al., 2016), we assume that the computational latency is proportional to the number of parameters. If the local LLM is comprised of $|\text{LLM}_{\text{Local}}|$ parameters and requires $2|\text{LLM}_{\text{Local}}|$ floating point operations (FLOPS) for forward propagation, the computational demand for each token in a reference processing length $|P|_{\text{ref}}$ is calculated as $\frac{2|\text{LLM}_{\text{Local}}|}{|P|_{\text{ref}}}$ FLOPS. The latency calculation considers the theoretical peak performance of the local device’s GPU, $TF_{\text{peak}}$ FLOPS. The computational latency can be computed as:
(13) $\small\psi_{\text{Local}}=\frac{2|\text{LLM}_{\text{Local}}|(|P|+|R|)}{|P|_{\text{ref}}\times TF_{\text{peak}}}.$
3.6.2. Cloud Interaction Latency $\psi_{\text{Cloud}}$
@@
The interaction time with the cloud LLM encompasses the entire process from sending the prompt to receiving the response. This includes uploading the prompt and modalities, inference by the cloud LLM, and downloading the response. We directly recorded the interaction time with the cloud LLM for different sets of uploaded modalities $\mathcal{M}$ during the creation of our M4A1 dataset.
3.7. Usage Cost Model
@@
During the inference process, LLMs incur computational costs. From the user’s perspective, for the local LLM, this manifests as energy consumption, while for the cloud LLM, it translates into the service fee paid to the provider.
3.7.1. Local LLM Usage Cost $\phi_{\text{Local}}$
@@
The energy consumption of the local LLM is influenced by inference time and the power consumption of the computing hardware, which are in turn affected by the local device’s computational capacity, the sizes of the prompt and response, and the size of the LLM, all encapsulated within the inference time $\psi_{\text{Local}}$. Let the maximum power consumption for a given local device hardware be $W_{\text{max}}$ Watts. By utilizing a normalized energy cost factor $\kappa$ (capturing e.g., local electricity rates and device wear-and-tear), after appropriate unit conversion, the local LLM usage cost (in USD) is defined as:
(14) $\phi_{\text{Local}}=\psi_{\text{Local}}\times W_{\text{max}}\times\kappa.$
3.7.2. Cloud LLM Usage Function $\phi_{\text{Cloud}}$
@@
For cloud-based LLMs, in addition to the inference costs of prompts and responses, the inference costs of data modalities must also be considered. These costs are typically available on the service provider’s website. Taking GPT-4o as an example, the prompt cost rate is $\varphi_{P}=\$0.005/1K$ tokens, the response cost rate is $\varphi_{R}=\$0.015/1K$ tokens, and the data modality costs are proportional to their size and resolution at a rate $\varphi_{m}$ for modality $m\in\mathcal{M}$. Therefore, the usage cost of the cloud LLM can be defined as follows:
(15) $\phi_{\text{Cloud}}=\varphi_{P}\cdot|P|+\varphi_{R}\cdot|R|+\sum_{m\in\mathcal{M}}\varphi_{m}\cdot\big{|}D^{(m)}\big{|}.$
Summary of Reward Function. Revisiting our initially defined reward function in Eq. (5), we substitute the direct response score $S_{R_{t}}$ with the estimated response score $S_{R_{t}}^{\prime}$ defined in Eq. (11) to achieve a more precise representation. The association metric is quantified through Eq. (12). Both the latency formulations ($\psi_{\text{Local}}$ and $\psi_{\text{Cloud}}$) and usage cost formulations ($\phi_{\text{Local}}$ and $\phi_{\text{Cloud}}$) are incorporated as $\psi_{a_{t}}$ and $\phi_{a_{t}}$ in the reward function, respectively.
4. Experimental Results
4.1. M4A1 Dataset
@@
To evaluate performance in practical settings, we develop M4A1, the first multi-modal, multi-task, multi-dialogue, and multi-LLM dataset, as shown in Fig. 3. M4A1 consists of 21,014 dialogues, each meticulously compiled by randomly and sequentially choosing tasks, prompts, modalities, and candidate LLMs to ensure comprehensive coverage across different configurations. The key characteristics of M4A1 can be summarized as follows:
@@
  • Four LLMs are a local LLM (Phi-3-mini), a cloud LLM (GPT-4o), a reference LLM (GPT-4o), and a judge LLM (GPT-4o).
@@
  • Four Metrics include response score (Sec. 3.4), association (Sec. 3.5), latency (Sec. 3.6), and usage cost (Sec. 3.7).
Figure 3. Illustration of M4A1 Dataset.
4.2. Experiment Setup
@@
Training Setup. The M4A1 dataset is split into a training set and a test set in an 8:2 ratio. The training set is used to train RLs while the test set is used to evaluate the performance at inference. When implementing our approach, we consider several popular discrete RL algorithms, including PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), and A2C (Mnih et al., 2016). We consider these methods with and without resource constraints (corresponding to cases with $\lambda>0$ and $\lambda=0$ in Eq. (10)), and denote RC-PPO, RC-DQN, and RC-A2C as the methods with the resource constraints. The policies are parameterized by multilayer perceptrons (MLP), with a total of 30,000 time steps set for training. For each experiment, we repeat the process three times and report the standard deviation as the error bar. We implement the TMO system and the baselines in PyTorch and conducted the experiments on an NVIDIA A100 GPU with 40 GB of memory.
Baselines. We compare the RL policies with three naive baselines: random (with random LLM and modality selection), local (using only the local LLM), and cloud (using only the cloud LLM, but with random modality selection). Inspired by recent advances in LLM-as-Agent (Xi et al., 2023; Liu et al., 2023), we substitute the RL agent with an LLM agent as TMO’s decision engine for the selection of LLMs and modalities. Specifically, current states and possible actions are translated into natural language descriptions as contextual input for the LLM agents, which then output an action in response to this prompt, following an RL-like procedure. The LLM agents include Phi-3-mini, Phi-3.5-mini, LLaMA-3.2-3B, LLaMA-3.1-8B, Mistral-7B-v0.3, FLAN-T5-large, FLAN-T5-xl, Gemma-2-2B, Gemma-2-9B, GPT-3.5-turbo, GPT-4o-mini, GPT-4o, OpenAI o1-mini, OpenAI o1, and OpenAI o3-mini. We reiterate that none of the existing works can be directly applied to our setting, which features multi-modal, multi-task, and multi-dialogue characteristics. Therefore, we reimplement AIwRG (He et al., 2024), an active inference method, and PerLLM (Yang et al., 2024), a bandit-based approach using upper confidence bound (UCB) for exploration-exploitation trade-off, integrating the settings introduced in this paper to adapt them to our research problem.
Default Settings. The default experimental settings are as follows unless otherwise stated; we also study the effects of various system parameters. Specifically, we use RC-A2C as the RL policy, with $\alpha=1$, $\beta_{\Lambda}=1/3$, $\beta_{\psi}=1/3$, and $\beta_{\phi}=1/3$ for the reward function weights, a 30 second latency constraint, a 0.05 USD usage cost constraint, and Jetson TX2 as the device type. These settings serve as the baseline from which variations are systematically introduced to assess their effects. Additionally, we set the time span $\tau=5$, the penalty coefficient $\lambda=10$, $k=5$ nearest neighbors for estimating the response score, and $\kappa=4.63\times 10^{-8}$ (Joules per USD) based on the average electricity price in the US  [^1] https://www.energybot.com/electricity-rates/.
Table 1. Main Result - TMO with RC-A2C optimally balances local/cloud LLMs and adaptively selects modalities, outperforming in response quality, latency, and costs without constraint violations. The highlighted row indicates the highest overall reward without constraint violations.
@@
Method Response Latency Usage Cost Overall Local Cloud - Num. Selected Modalities Constraint Violation ($\downarrow$)
@@
LLM-as-Agent (Xi et al., 2023; Liu et al., 2023) for Inference Offloading (Ignored LLM Agent’s Latency and Usage Cost)
@@
Comparison with SOTA Exploration-Decision Baselines
AIwRG (He et al., 2024) 1.02 $\pm$0.05 23.73 $\pm$0.98 44.01 $\pm$3.50 0.69 $\pm$0.06 113 244 0 0 3852 4.68 $\pm$0.18 9.26 $\pm$1.83
PerLLM (Yang et al., 2024) 0.97 $\pm$0.06 21.12 $\pm$0.08 1.60 $\pm$0.01 0.66 $\pm$0.06 281 3922 0 0 0 0.00 $\pm$0.00 0.00 $\pm$0.00
TMO (PPO) 1.02 $\pm$0.00 17.89 $\pm$0.82 22.48 $\pm$1.48 0.76 $\pm$0.01 142 40 2499 1355 121 0.70 $\pm$0.77 1.66 $\pm$2.35
TMO (DQN) 1.02 $\pm$0.08 19.57 $\pm$2.18 26.23 $\pm$6.36 0.75 $\pm$0.05 0 5 1820 2367 0 0.81 $\pm$0.57 0.00 $\pm$0.00
TMO (A2C) 1.05 $\pm$0.03 18.81 $\pm$3.44 26.44 $\pm$13.05 0.79 $\pm$0.07 98 0 2753 210 1168 1.48 $\pm$2.09 2.38 $\pm$3.36
TMO (RC-PPO) 1.03 $\pm$0.03 18.33 $\pm$1.87 22.43 $\pm$3.57 0.77 $\pm$0.04 90 105 2599 1309 110 0.45 $\pm$0.64 0.00 $\pm$0.00
TMO (RC-DQN) 1.05 $\pm$0.05 18.87 $\pm$0.81 23.29 $\pm$3.59 0.78 $\pm$0.06 46 174 2249 1636 85 0.53 $\pm$0.33 0.42 $\pm$0.60
TMO (RC-A2C) 1.09 $\pm$0.08 16.63 $\pm$0.67 17.97 $\pm$1.18 0.85 $\pm$0.07 74 12 3871 225 6 0.00 $\pm$0.00 0.00 $\pm$0.00
Our TMO System - Ablation Study (Using RC-A2C as Backbone)
w/o LLM Sel. 0.85 $\pm$0.00 9.25 $\pm$1.02 11.67 $\pm$3.14 0.72 $\pm$0.02 2128 6 1228 830 1 0.00 $\pm$0.00 0.00 $\pm$0.00
@@
4.3. Main Experimental Results